# Detection of disk-jet co-precession in a tidal disruption event

### v1

@@ 4 Radio long-term evolution
where the best-fitting break time is $x_{\rm break}=167.9\pm 6.3$ day. The slope during the rising phase is $a_{1}=1.8\pm 0.4$, and the slope during the decaying phase is $a_{2}=-1.2\pm 0.2$. The normalization parameter is $b=1.8\pm 0.1\times 10^{38}$$\mathrm{erg\,s^{-1}}$. We then select three additional epochs that show the least deviation from the broken power-law function and fitted the epochs, for which only C-band data is available, using the obtained Gaussian priors for $\log{F_{1}}$ and $p$. The results are shown in Fig. S5.
@@
5 Nature of the 2024 re-brightening
[…]
In the two latest NTT spectra (2024-08-27 and 2024-12-24, at +239 and +358 days, respectively), we detected high-ionization coronal lines (CLs). These include Fe xiv $\lambda 5303$, Fe vii $\lambda\lambda 5720,6087$, and Fe x $\lambda 6374$. There are indications of CLs in the first NTT spectrum taken before the seasonal gap, although they appear significantly weaker. CLs are high-ionization lines (e.g., of Iron, Neon, Argon, Sulfur) that originate from the photoionization or collisional ionization of a clumpy interstellar medium or other pre-existing material. Although typically associated with AGN, strong CLs have been detected in the optical spectra of some galaxies that show little to no evidence of AGN activity (?, ?). Many of these CLs have ionization potentials of $\geq 100$ eV, requiring a strong extreme UV (EUV) and/or soft X-ray ionizing continuum. It has long been thought that such extreme coronal line emission events (ECLEs) are powered by TDEs, and recent studies support this scenario (?, ?). Interestingly, a few optical-UV TDE candidates (?, ?) out of the $\sim 50$ strong candidates discovered to date have exhibited late ($\geq 200$ days post-peak) X-ray brightening accompanied by the emergence of iron coronal lines. However, given that the strength of these CLs increased as the high-energy emission decreased, combined with the weak accretion activity detected in 2020 in AT2020afhd, these lines are likely associated with the interaction between outflows and pre-existing material, or with debris containing iron that only reached the optical emission site at later times. Further investigation with detailed modeling is needed to distinguish the two scenarios and will be addressed in a follow-up paper.
@@ 6 Host galaxy properties
To constrain the host galaxy properties and estimate its contribution to the UVOT and VT detection, we used the Python package CIGALE for SED fitting. We gathered archival observations covering wavelengths from the UV to the mid-IR. For the mid-IR, we utilized W1 and W2 magnitudes from the AllWISE Source Catalog. For the near-IR, we obtained J, H, and Ks magnitudes from the 2MASS All-Sky Point Source Catalog. For the optical bands, we calculated the mean g, r, i, z, and y magnitudes from the PS1 Detection table during the inactive period (MJD 55800–55880) of the host galaxy. For the UV, we extracted GALEX NUV and FUV magnitudes from the All-sky Imaging Survey in the GALEX DR5 catalog.
@@ 7 Black hole mass estimation
We independently measured the black hole (BH) mass using the single-epoch method and the M-sigma relation. These measurements were based on the medium-resolution spectrum obtained from DESI taken prior to the TDE.
Before deriving the properties of the emission lines, it was crucial to accurately subtract the host galaxy’s continuum and absorption features. To achieve this, we used the Penalized PiXel Fitting software (pPXF), originally developed by ref. (?) and further refined by ref. (?). The pPXF incorporates stellar templates from the MILES Library (?). This library contains 985 well-calibrated stellar spectra covering a wavelength range of 3525–7500 Å, with a spectral resolution characterized by a full width at half maximum (FWHM) of 2.51 Å and a velocity dispersion ($\sigma$) of approximately 64 km s-1.
@@ 8 Model setup and constraints
In TDEs, the nascent accretion disk and the jet, which is perpendicular to the disk, are highly likely to be misaligned with the spinning BH’s equatorial plane, resulting in LT precession due to spacetime frame-dragging. The disk’s structure and dynamics depend on the accretion rate: at super-Eddington rates, the disk is geometrically thick with the semi-thickness, $H/R\geq\alpha$, where $H$ and $R$ is the disk’s height and radius, respectively, and $\alpha$ is the viscosity parameter (?). Under these conditions, the disk undergoes rigid-body precession, moving as a single unit. As the accretion rate decreases to sub-Eddington levels, the disk becomes thin, $H/R<\alpha$, leading to significant warping and precession in the inner region while the outer region remains stable and retains its initial inclination. After this, the Bardeen-Petterson (BP, (?)) effect begins to dampen the LT precession, resulting in an alignment between the disk and the BH’s angular momentum.
For the TDE disk, we adopted a commonly used model in TDE studies as described in (?, ?). Following the disruption of a solar-type star by a SMBH with $M_{\bullet}\sim 10^{6.7\pm 0.5}\rm M_{\odot}$, the accretion rate is expected to exceed the Eddington limit during the first year. In this super-Eddington phase, the disk is geometrically thick, with a scaleheight ratio of $H/R\sim 1$ (?). Thus, in this study, we mainly focused on the super-Eddington phase, where the disk was resistant to Bardeen-Petterson warps (?, ?).
@@ 8.1 X-ray variations and black hole spin estimation
To test whether the quasi-periodic variations in X-rays could be attributed to disk LT precession, we investigated the X-ray flux as a function of the angle $\psi(t)=\cos^{-1}(\hat{r}_{\rm{obs}}\cdot\hat{l}_{\rm{disk}}(t))$, which represents the angle between the observer’s line of sight (the direction is denoted by the unit vector $\hat{r}_{\rm{obs}}$) and the disk’s angular momentum (the direction at time $t$ is denoted by the unit vector $\hat{l}_{\rm{disk}}(t)$). $\hat{r}_{\rm{obs}}$ is parameterized by the viewing angle $\theta_{\rm obs}$ (with respect to the BH spin axis $Z$) and $\phi_{\rm obs}$ as shown in Fig.5. The disk precessed with an angle $\theta_{\rm i}$ around the BH spin axis. Consequently, $\hat{l}_{\rm{disk}}(t)$, and thus $\psi(t)$, evolved over time due to LT precession (see Fig. 5).
@@
where $r=R/R_{\rm g}$, with $R_{\rm{g}}=GM_{\bullet}/c^{2}$ being the gravitational radius of the BH, $f_{\rm g}=1-(R_{\rm ms}/R)^{1/2}$, with $R_{\rm ms}$ being the innermost stable circular orbit (ISCO), and $\epsilon=1-E_{\rm ms}$ is the radiative efficiency of disk, with $E_{\rm ms}$ being the specific energy at $R_{\rm ms}$ (?). $R_{\rm ms}$ and $E_{\rm ms}$ are functions of the BH spin parameter $a_{\bullet}$, which can be constrained using the LT precession period as follows (see also Fig. S8b). The parameter $\dot{m}$ is the disk mass accretion rate $\dot{M}$ normalized to the Eddington accretion rate, $\dot{M}_{\rm Edd}=L_{\rm Edd}/c^{2}/\epsilon$. Here, $L_{\rm Edd}=1.26\times 10^{44}M_{\bullet,6}\ \rm erg\ s^{-1}$ is the Eddington luminosity, with $M_{\bullet,6}$ representing the BH mass scaled by $10^{6}M_{\odot}$. In the super-Eddington phase, the disk thickness increases with the radius, reaching a maximum $H/R\sim 1$ near the disk outer radius $R_{\rm out}$ (?). In our model, we set $R_{\rm out}$ to the circularization radius $R_{\rm cir}$ of the debris, such that $R_{\rm out}=R_{\rm cir}\simeq 2R_{\rm tde}$, assuming the impact parameter $\beta=R_{\rm tde}/R_{\rm p}\simeq 1$ (?). Here, $R_{\rm{tde}}\simeq(M_{\bullet}/M_{*})^{1/3}R_{*}$ represents the tidal disruption radius, where $M_{*}$ and $R_{*}$ are the mass and radius of the disrupted star, and $R_{\rm p}$ is the pericentre of its orbit. Therefore, at large $\psi$, the X-ray emission from the inner disk is obscured by the outer, thicker portion of the disk, resulting in the observed sharp dips in the X-ray lightcurve. The X-ray variations are attributed to changes in the disk’s projected area and periodic obscuration by the outer disk. For the same reason, the disk temperature should also modulate over the rigid-body precession period (observe the high-temperature inner part of the disk/X-ray peak when viewed at small $\psi$, and the low-temperature outer part/X-ray dip at the phases of large $\psi$), which is consistent with the observations in AT2020afhd (Fig. 2a). This supports the precession of an extended disk rather than a narrow ring (?).
@@
Using the MCMC approach (?), we fitted the 0.1–2 keV X-ray lightcurve, observed between August 3 and October 21, 2024, where the X-ray quasi-periodic variations were clearly detected, with the proposed model. The model assumed a disk thickness ratio of $H/R\sim 1$ and an accretion rate of $\dot{m}\sim 1$. Flat priors were applied, including $0^{\circ}<\theta_{\rm i}<30^{\circ}$ for the disk precession angle, $0^{\circ}<\theta_{\rm obs}<90^{\circ}$ for the observer’s inclination/viewing angle, $0^{\circ}<\phi_{\rm obs}<360^{\circ}$ for the azimuthal angle between the disk and the line of sight, and the variation amplitude constrained to exceed one order of magnitude. We initialized 30 walkers to explore the parameter space, each taking 10,000 steps to solve for the maximum likelihood solution. The best-fitting parameters were $\theta_{\rm obs}\sim 38.4^{+0.5\circ}_{-0.6}$, $\theta_{\rm i}\sim 14.5^{\circ}\pm 0.5^{\circ}$, and $\phi_{\rm obs}\sim 188.4^{\circ}\pm 1.5^{\circ}$, as shown in Fig. S8a.
@@
$T_{\rm prec}=\frac{2\pi}{\Omega_{\rm p}}=\frac{\pi GM_{\bullet}r_{\rm ms}^{3}}{a_{\bullet}c^{3}}\frac{1-x_{\rm out}^{-1/2-p}}{x_{\rm out}^{5/2-p}-1}\frac{5/2-p}{p+1/2},$ (S4)
where $\Omega_{\rm p}$ is the precession frequency, $x_{\rm out}=R_{\rm out}/R_{\rm ms}$ and $r_{\rm ms}=R_{\rm ms}/R_{\rm g}$. Using power-law indices for a surface density profile of $p=0,3/5,3/4$ (?), a SMBH mass of $M_{\bullet}=10^{6.7\pm 0.5}~M_{\odot}$ and assuming the disruption of a solar-like star, we found that $a_{\bullet}$ fell within the range of (0.460.14)-(0.46-0.14) or 0.110.350.11-0.35 (Fig. S8). A negative spin parameter would have resulted in a larger innermost stable circular orbit, potentially suppressing the high-energy X-ray emission from the accretion disk or even hindering the occurrence of the TDE (i.e., $R_{\rm ms}>R_{\rm tde}$). Therefore, a positive spin parameter was favored.
For AT2020afhd, the X-ray quasi-periodic variations disappeared after approximately 300 days. This result is consistent with the expected alignment timescale, $t_{\rm align}\sim(1/\alpha)(H/R)^{2}(\Omega/\Omega_{\rm p}^{2})$, assuming $\alpha\sim 0.3$ and the viscous dissipation as the main alignment mechanism (?, ?). Here, $\Omega=c^{3}/(GM_{\bullet})(1/(r^{3/2}+a_{\bullet}))$ represents the disk angular velocity. Another possibility is that the disk became thinner as the accretion rate continued to decline, causing $H/R$ to drop below $\alpha$. In this state, the rigid-body precession ceased, and the disk rapidly aligned with the BH’s spin (?, ?). For $M_{\bullet}\sim 10^{6.7}M_{\odot}$ and $\alpha\sim 0.3$, the rigid body precession should last $\sim 300$ days before the disk becomes geometrically thin (?).
@@ 8.2 Radio variations
For an observer at $\psi$ relative to the jet axis (as shown in Fig. 5), the observed flux could be expressed as (?, ?):
@@
where $a_{\rm off}(\psi,\Gamma)={\mathcal{D}}_{\rm off}/{\mathcal{D}}_{\rm on}=(1-\beta_{\rm j})(1-\beta_{\rm j}\cos\psi)$ is the ratio of the on-beam to the off-beam Doppler factor, and $\beta_{\rm j}=\sqrt{1-1/\Gamma^{2}}$. For a TDE, the long-term lightcurve is defined by the fallback accretion rate, which is proportional to $t^{-5/3}$, and therefore have $F(\psi,t)=a_{\rm off}^{7/3}(\psi,\Gamma)F(0,t)$. The radio variations exhibited a peak-to-dip ratio of $F(\psi_{\rm peak})/F(\psi_{\rm dip})$ exceeding $4$. Using $\theta_{\rm obs}\sim 38.4^{\circ}$ and $\theta_{\rm i}\sim 14.5^{\circ}$, as suggested by the disk precession model, we determined $\psi_{\rm peak}=\theta_{\rm obs}-\theta_{\rm i}\simeq 23.9^{\circ}$ and $\psi_{\rm dip}=\theta_{\rm obs}+\theta_{\rm i}\simeq 52.9^{\circ}$. From this, we constrained the jet’s Lorentz factor to $\Gamma\leq 1.6$.
@@
Compared to jetted TDEs such as Swift J1644+57, the derived high BH spin ($a_{\bullet}\geq 0.6$) (?, ?, ?) and strong magnetic field ($\sim 10^{6}$ G) (?, ?) account for its powerful jet, with internal dissipation primarily emitting in X-rays. In contrast, the lower BH spin and weaker magnetic field in AT2020afhd explain its weaker jet, where internal dissipation is predominantly observed in the radio band.
9 Other potential mechanisms driving X-ray and radio co-variances
Disk radiation pressure instabilities and the orbital motion of binary SMBH systems may also produce comparable periodicity in X-ray and radio emissions. For example, the microquasar GRS 1915+105 exhibited simultaneous periodic modulations in X-ray and radio bands on timescales of 20–50 minutes (?, ?, ?). This periodic flaring activity has been attributed to radiation pressure instabilities, where material in the inner region of an optically thick accretion disk is rapidly depleted and replenished (?, ?). During this process, part of the inner disk is ejected to form a jet, as indicated by a rise in the radio flare coinciding with a dip in the X-ray flare (?). However, the X-ray and radio variations in GRS 1915+105 are considerably more complex, resulting in diverse correlations between the two-band lightcurves (?). Overall, the lack of simulations and theoretical models on the short-term evolution of jets during radiation-pressure instabilities makes it challenging to further test this scenario. While in binary SMBH systems, the variability period and profile caused by orbital motion are generally not expected to evolve on observational timescales, making this inconsistent with the behaviors observed in AT2020afhd.
Figure S1: PanSTARRS-$g$ band image centered at the position of the galaxy LEDA 145386. The colored crosses mark the centers of the multi-wavelength counterparts of AT2020afhd. The purple and green dashed circles, with radii of 0.5” and 1.2”, respectively, indicate the positional uncertainties. The positional uncertainty for VLBA is too small to display.
@@ 9 Other potential mechanisms driving X-ray and radio co-variances
Figure S3: Illustration of four selected X-ray spectra. The XRT (ObsID: 00016492036), NICER (ObsID: 6705010126), and PN (ObsID: 0953010201) spectra were obtained within 1.2 days after the TDE, while the eROSITA spectrum was taken prior to the TDE.
Figure S4: a, The CDF of the normalized LSP noise power from the observed data (blue points) compared to the expected CDF of an exponential distribution (red curve). b, The PDF of the normalized LSP noise power from the observed data (blue histogram) compared to the expected PDF of an exponential distribution (red curve). c, K-S test results from simulations. The orange histogram shows the distribution of K-S statistic values for simulated white noise, with the $1\sigma$ range shaded in grey. The red line represents the K-S statistic value for the observed LSP noise power. d, A-D goodness-of-fit test results from simulations. The orange histogram shows the distribution of A-D statistic values for simulated white noise, with the $1\sigma$ range shaded in grey. The red line represents the A-D statistic value for the observed LSP noise power.
@@
Figure S6: a, ZTF and VT photometries of AT2020afhd. The ZTF data were corrected for extinction, while the VT data were corrected for Galactic extinction and had the host contribution subtracted. The dashed line indicates the magnitude level of the detection in 2020. b, Host-galaxy SED and the best-fitting model.
[…]
Figure S8: a, Contour plots showing the best-fitting parameters for our proposed disk precession model. b, Disk precession period vs BH spin when adopting various BH masses and surface density profile of the accretion disk. The gray and the blue shades correspond to a BH mass of $10^{6.7\pm 0.5}~M_{\odot}$ and a period of $19.6\pm 1.5$ days. The overlap of the two shades constrains the BH spin to ranges of $0.46<a_{\bullet}<-0.14$ and $0.11<a_{\bullet}<0.35$.
Table S1: Summary of the spectroscopic observations
4d P60/SEDM - 2024-01-13T06:00:32 $2700$
43d FLOYDS-S - 2024-02-13T10:36:20 $1800$
[…]
Note. Spectra from P60 and FLOYDS were downloaded directly from the Transient Name Server******https://www.wis-tns.org/object/2020afhd. The epochs correspond to the times offset from MJD 60310. Spectra with a SNR below 10 are excluded from the table, as they are not utilized in the subsequent analysis.

### v2

@@ 3 Radio long-term evolution
where the best-fitting break time is $x_{\rm break}=167.9\pm 6.3$ day. The slope during the rising phase is $a_{1}=1.8\pm 0.4$, and the slope during the decaying phase is $a_{2}=-1.2\pm 0.2$. The normalization parameter is $b=1.8\pm 0.1\times 10^{38}$$\mathrm{erg~s^{-1}}$. We then select three additional epochs that show the least deviation from the broken power-law function and fitted the epochs, for which only C-band data is available, using the obtained Gaussian priors for $\log{F_{1}}$ and $p$. The results are shown in fig. S4.
@@ 4 Host galaxy properties
To constrain the host galaxy properties and estimate its contribution to the UVOT and VT detection, we used the Python package CIGALE for SED fitting. We gathered archival observations covering wavelengths from the UV to the mid-IR. For the mid-IR, we utilized W1 and W2 magnitudes from the AllWISE Source Catalog. For the near-IR, we obtained J, H, and Ks magnitudes from the 2MASS All-Sky Point Source Catalog. For the optical bands, we calculated the mean g, r, i, z, and y magnitudes from the PS1 Detection table during the inactive period (MJD 55800–55880) of the host galaxy. For the UV, we extracted the NUV and FUV magnitudes from the All-sky Imaging Survey in the GALEX DR5 catalog.
@@ 5 Black hole mass estimation
We independently measured the black hole (BH) mass using the single-epoch method and the M-sigma relation. These measurements were based on the medium-resolution DESI spectrum taken prior to the TDE.
Before deriving the properties of the emission lines, it was crucial to accurately subtract the host galaxy’s continuum and absorption features. To achieve this, we used the Penalized PiXel Fitting software (pPXF), originally developed by (?) and further refined by (?). The pPXF incorporates stellar templates from the MILES Library (?). This library contains 985 well-calibrated stellar spectra covering a wavelength range of 3525–7500 Å, with a spectral resolution characterized by a FWHM of 2.51 Å and a velocity dispersion ($\sigma$) of approximately 64 km s-1.
@@ 6 Model setup and constraints
In TDEs, the nascent accretion disk and the associated jet are highly likely to be misaligned with the spinning BH’s equatorial plane, resulting in LT precession due to spacetime frame-dragging. Following the disruption of a solar-type star by a SMBH with $M_{\bullet}\sim 10^{6.7\pm 0.5}~M_{\odot}$, the accretion rate is expected to exceed the Eddington limit during the first year. During this phase, the disk is geometrically thick with the semi-thickness, $H/R\geq\alpha$, where $H$ and $R$ is the disk’s height and radius, respectively, and $\alpha$ is the viscosity parameter (?). Under these conditions, the disk undergoes rigid-body precession, moving as a single unit.
In the super-Eddington accretion regime of TDEs, the thermal luminosity is expected to remain roughly constant as the accretion rate declines (?). The two plateaus observed in the X-ray lightcurve of AT2020afhd prior to day 300 likely indicate periods of sustained super-Eddington or Eddington-limited accretion. Overall, we adopt a commonly used TDE disk model described in (?, ?), where the disk remains geometrically thick with $H/R\sim 1$. Throughout, we assume that the jet is launched perpendicular to the plane of the disk.
@@ 6.1 X-ray variations and black hole spin estimation
To test whether the quasi-periodic variations in X-rays could be attributed to disk LT precession, we investigated the X-ray flux as a function of the angle $\psi(t)=\cos^{-1}(\hat{r}_{\rm{obs}}\cdot\hat{l}_{\rm{disk}}(t))$, which represents the angle between the observer’s line of sight (the direction is denoted by the unit vector $\hat{r}_{\rm{obs}}$) and the disk’s angular momentum (the direction at time $t$ is denoted by the unit vector $\hat{l}_{\rm{disk}}(t)$). $\hat{r}_{\rm{obs}}$ is parameterized by the viewing angle $\theta_{\rm obs}$ (with respect to the BH spin axis $Z$) and $\phi_{\rm obs}$ as shown in Fig. 5A. The disk precesses with an angle $\theta_{\rm i}$ around the BH spin axis. Consequently, $\hat{l}_{\rm{disk}}(t)$, and thus $\psi(t)$, evolved over time due to LT precession.
@@
where $r=R/R_{\rm g}$, with $R_{\rm{g}}=GM_{\bullet}/c^{2}$ being the gravitational radius of the BH, $f_{\rm g}=1-(R_{\rm ms}/R)^{1/2}$, with $R_{\rm ms}$ being the innermost stable circular orbit (ISCO), and $\epsilon=1-E_{\rm ms}$ is the radiative efficiency of disk, with $E_{\rm ms}$ being the specific energy at $R_{\rm ms}$ (?). $R_{\rm ms}$ and $E_{\rm ms}$ are functions of the BH spin parameter $a_{\bullet}$, which can be constrained using the LT precession period (see also Fig. 6A). The parameter $\dot{m}$ is the disk mass accretion rate $\dot{M}$ normalized to the Eddington accretion rate, $\dot{M}_{\rm Edd}=L_{\rm Edd}/c^{2}/\epsilon$. Here, $L_{\rm Edd}=1.26\times 10^{44}M_{\bullet,6}\ \rm erg\ s^{-1}$ is the Eddington luminosity, with $M_{\bullet,6}$ representing the BH mass scaled by $10^{6}M_{\odot}$. In the super-Eddington phase, the disk thickness increases with the radius, reaching a maximum $H/R\sim 1$ near the disk outer radius $R_{\rm out}$ (?). In our model, we set $R_{\rm out}$ to the circularization radius $R_{\rm cir}$ of the debris, such that $R_{\rm out}=R_{\rm cir}\simeq 2R_{\rm tde}$, assuming the impact parameter $\beta=R_{\rm tde}/R_{\rm p}\simeq 1$ (?). Here, $R_{\rm{tde}}\simeq(M_{\bullet}/M_{*})^{1/3}R_{*}$ represents the tidal disruption radius, where $M_{*}$ and $R_{*}$ are the mass and radius of the disrupted star, and $R_{\rm p}$ is the pericentre of its orbit. Therefore, at large $\psi$, the X-ray emission from the inner disk is obscured by the outer, thicker portion of the disk, resulting in the observed sharp dips in the X-ray lightcurve. The X-ray variations are attributed to changes in the disk’s projected area and periodic obscuration by the outer disk. For the same reason, the disk temperature should also modulate over the rigid-body precession period (observe the high-temperature inner part of the disk/X-ray peak when viewed at small $\psi$, and the low-temperature outer part/X-ray dip at large $\psi$), which is consistent with the observations in AT2020afhd (Fig. 2A). This supports the precession of an extended disk rather than a narrow ring (?).
@@
Using the MCMC approach (?), we fitted the 0.1–2 keV X-ray lightcurve, observed between August 3 and October 21, 2024, where the X-ray quasi-periodic variations were clearly detected, with the proposed model. The model assumes a disk thickness ratio of $H/R\sim 1$ and an accretion rate of $\dot{m}\sim 1$. Flat priors were applied, including $0^{\circ}<\theta_{\rm i}<30^{\circ}$ for the disk precession angle, $0^{\circ}<\theta_{\rm obs}<90^{\circ}$ for the observer’s inclination/viewing angle, $0^{\circ}<\phi_{\rm obs}<360^{\circ}$ for the azimuthal angle between the disk and the line of sight, and the variation amplitude constrained to exceed one order of magnitude. We initialized 30 walkers to explore the parameter space, each taking 10,000 steps to solve for the maximum likelihood solution. The best-fitting parameters were $\theta_{\rm obs}\sim 38.4^{+0.5\circ}_{-0.6}$, $\theta_{\rm i}\sim 14.5^{\circ}\pm 0.5^{\circ}$, and $\phi_{\rm obs}\sim 188.4^{\circ}\pm 1.5^{\circ}$, as shown in Fig. 6B.
@@
$T_{\rm prec}=\frac{2\pi}{\Omega_{\rm p}}=\frac{\pi GM_{\bullet}r_{\rm ms}^{3}}{a_{\bullet}c^{3}}/(\frac{1-x_{\rm out}^{-1/2-p}}{x_{\rm out}^{5/2-p}-1}\frac{5/2-p}{p+1/2}),$ (S4)
where $\Omega_{\rm p}$ is the precession frequency, $x_{\rm out}=R_{\rm out}/R_{\rm ms}$ and $r_{\rm ms}=R_{\rm ms}/R_{\rm g}$. Using power-law indices for a surface density profile of $p=0,3/5,3/4$ (?), a SMBH mass of $M_{\bullet}=10^{6.7\pm 0.5}~M_{\odot}$ and assuming the disruption of a solar-like star, we found that $a_{\bullet}$ fell within the range of (0.460.14)-(0.46-0.14) or 0.110.350.11-0.35 (Fig. 6A). A negative spin parameter would have resulted in a larger innermost stable circular orbit, potentially suppressing the high-energy X-ray emission from the accretion disk. Therefore, a positive spin parameter was favored.
Approximately 300 days after the re-brightening, the 19.6-day quasi-periodic variations disappeared, and the correlation between the X-ray and radio emissions broke down, suggesting that the angular momenta of the accretion disk and the SMBH may have aligned by that time. Assuming a 300-day alignment timescale and adopting the system parameters derived above, one can place constraints on the disk viscosity parameter, $\alpha$. To this end, we consider two possible alignment mechanisms: viscous dissipation and disk cooling (?, ?). In both scenarios, $\alpha$ is allowed to vary freely within the range 0.01 to 1. When viscous dissipation dominates the alignment (?, ?), the alignment timescale can be expressed as:
$t_{\rm align}\sim(1/\alpha)(H/R)^{2}(\Omega/\Omega_{\rm p}^{2}),$ (S5)
where $\Omega=c^{3}/(GM_{\bullet})(1/(r^{3/2}+a_{\bullet}))$ represents the disk angular velocity. Using this relation yields $\alpha\sim 0.6$. Notably, the choice of $a_{\bullet}$ does not substantially affect this result. Alternatively, if disk cooling drives the alignment (?, ?), the timescale can be estimated as:
$t_{\rm thin}=t_{\min}\alpha^{-3/5}\left[5\,f_{\rm g}\frac{M_{*}}{\dot{M}_{\rm Edd}t_{\min}}\frac{1}{r}\right]^{3/5},$ (S6)
where $t_{\min}=41M_{\bullet,6}^{1/2}(M_{*}/M_{\odot})^{-1}(R_{*}/R_{\odot})^{3/2}$ days represents the timescale for the disrupted material first coming back to pericentre. Adopting $a_{\bullet}\sim 0.11-0.35$ yields $\alpha\sim 0.2-0.4$. In principle, the actual alignment timescale could exceed 300 days if the accretion disk had begun precessing prior to the observed re-brightening. In such a case, smaller values of $\alpha$ would be required under both alignment mechanisms.
@@ 6.2 Radio variations
For an observer at $\psi$ relative to the jet axis (as shown in Fig. 5A), the observed flux could be expressed as (?, ?):
@@
where $a_{\rm off}(\psi,\Gamma)={\mathcal{D}}_{\rm off}/{\mathcal{D}}_{\rm on}=(1-\beta_{\rm j})(1-\beta_{\rm j}\cos\psi)$ is the ratio of the on-beam to the off-beam Doppler factor, and $\beta_{\rm j}=\sqrt{1-1/\Gamma^{2}}$. For a TDE, the long-term lightcurve is defined by the fallback accretion rate, which is proportional to $t^{-5/3}$, and therefore have $F(\psi,t)=a_{\rm off}^{7/3}(\psi,\Gamma)F(0,t)$. The radio variations exhibited a peak-to-dip ratio, $F(\psi_{\rm peak})/F(\psi_{\rm dip})$, ranging from 1.9 to 4.3. Using $\theta_{\rm obs}\sim 38.4^{\circ}$ and $\theta_{\rm i}\sim 14.5^{\circ}$, as suggested by the disk precession model, we determined $\psi_{\rm peak}=\theta_{\rm obs}-\theta_{\rm i}\simeq 23.9^{\circ}$ and $\psi_{\rm dip}=\theta_{\rm obs}+\theta_{\rm i}\simeq 52.9^{\circ}$. From this, we constrained the jet’s Lorentz factor to $1.2\leq\Gamma\leq 1.6$.
@@
Figure S1: PanSTARRS-$g$ band image centered at the position of the galaxy LEDA 145386. The colored crosses mark the centers of the multi-wavelength counterparts of AT2020afhd. The purple and green dashed circles, with radii of 0.5′′ and 1.2′′, respectively, indicate the positional uncertainties. The positional uncertainty for VLBA is too small to display.
@@
Figure S3: Illustration of four selected X-ray spectra. (A to C) The XRT (ObsID: 00016492036), NICER (ObsID: 6705010126), and PN (ObsID: 0953010201) spectra were obtained within 1.2 days after the TDE, while (D) the eROSITA spectrum was taken prior to the TDE.
@@
Figure S5: Some optical properties of AT2020afhd. (A) ZTF and VT photometries. The ZTF data were corrected for extinction, while the VT data were corrected for Galactic extinction and had the host contribution subtracted. The dashed line indicates the magnitude level of the detection in 2020. (B) Host-galaxy SED and the best-fitting model.